You have two Windows Server 2012 R2 computers with the Hyper-V role installed. Both computers have two hard drives, one for the system volume and the other for data. One server, named HyperVTest, is going to be used mainly for testing and what-if scenarios, and its data drive is 250 GB. You estimate that you might have 8 or 10 VMs configured on HyperVTest with two or three running at the same time. Each test VM has disk requirements ranging from about 10 GB to 30 GB. The other server, named HyperVApp, runs in the data center with production VMs installed. Its data drive is 500 GB. You expect two VMs to run on HyperVApp, each needing about 150 GB to 200GB of disk space. Both are expected to run fairly disk-intensive applications. Given this environment, describe how you would configure the virtual disks for the VMs on both servers.

Please use complete sentences and be detailed in the answer.

The following that was answered is NOT correct: For HyperVTest server, in the worst case, 3 VMs will be active with maximum 30 GB disk requirement. Hence, space required = 3*30 = 90 GB in 3 equal virtual disks. For HyperVApp, 2 virtual disk should have 200 GB each. Total size of virtual disk = 350 GB
Answer #1

HyperVTest :
Data Drive : 250 GB
Configured VM's = 8-10
Concurrent running VM's = 2-3
Disk Requirement = 10-30 GB
As the Data Drive is of size 250 GB we can configure 3 30GB VMs, 4 25GB VMs and 3 20GB VMs.

HyperVApp :
Data Drive : 500 GB
Concurrent running VM's = 2
Disk Requirement = 150-200 GB
As the Data Drive is of size 500 GB we can have 1 200GB VM and 2 150GB VM configured.